An engine whestles at a frequency of 256Hz moves towards  a wall with a velocity of 50m/s. How many number of beats will be sounded to the engine?

Grade:12

1 Answers

devashish trivedi
15 years ago

Doppler's fomula is

           n' = n (v -vo)/(v-vs)

where n' = apprant frequency

           n = original frequency

         v = spped of sound in air that is 330 m/s

      vo = speed of observer

     vs = speed of the source

Here n = 256 hz vo = 0 vs = 50m/s pluging in the values we will get n = 497 hz

now with this frequency the sond will reflect from the wall so the source will be stationary

we will again use the same formula

                                                       n" = n (v -vo)/(v-vs)

now vo = 50m/s , vs = 0 and v =330m/s n" = frequency heard by the observer

now pluging in the values we will have

                                                    n" = 497(330 - 50)/(330 - 0)

                                                           = 497x 280/330

                                                          = 422hz so the no of beats heard by the observer is n" - n = 422 - 256=      166 beats per second